Accountancy

The profit of mahesh and company for the last 5 years were 

2005_35000, 2006_30, 000,2007_27500, 2008_25000, 2009_20000.The capital employed in the form is 500,000.you are required to calculate the goodwill at 3 years purchase of super profits. The normal rate of return on capital employed is 5% 
Answer #1

Average profit = 35000+30000+27500+25000+20000/5

=137500/5=27500

Normal profit=500000×5/100=25000

Super profit=Actual average profit_ normal profit 

=27500-25000=2500

Goodwill=2500×3=7500

answered by: Advika